9. Your investment advisor expects that on ABC S.A. you can earn 0,5PLN per share next year, 0,6PLN the year after that, and then the earnings grow by 8% each year. Dividend payout is 60% and the required rate of return is 15%. Price per share is 4 PLN. What would you do with this proposal: buy, sell or retain?
Answer #1

SEE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE. THANK YOU

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard BY112 BS Alignment Number Cells BT BU BV BW BX BY BZ DPS (60%) 105 106 107 108 109 110 STEP 1 year EPS 0.5 0.6 0.3 0.36 0.3888 0.648 0.6*(1+0.08) P2 = D3/ke-g) P2 0.3888/(0.15-0.08) P2 = 112 113 114 115 116 5.55 STEP 2 PV FACTOR AT 15% PV OF YEAR DPS+P2 DPS P2 0.5 6.15 0.8696 0.7561 0.43 4.65 118 119 120 121 122 123 14 1 r -1 | KE CAPM / UTILITY, SH (0.6+ 5.55) P0= 5.09 AS CURRENT PRICE IS 4, LESS THAN INTRINSIC VALUE (5.09) YOU SHOULD BUY THE STOCK AND YOU ALREADY OWN, RETAIN IT STEP 3 beta bond c M port future INDEX INTL CAP BUD LEASING PV, FV, ANNUITY DIR cleanYIELD bond stru WACC RES1 ex di erences: GW40 福 130% 17:49 15-01-2019